# In Situ AI Prototyping: Infusing Multimodal Prompts into Mobile Settings with MobileMaker

### v1

@@ VI-B2 NL revise enabled users to critically evaluate their own feedback in situ
With MobileMaker’s NL revision feature, participants could immediately “live” and experience their own feedback within the updated prototype, which often helped them rethink their initial feedback. P4 initially revised Img2VideoIdeas to “Add a script for each video idea”, but after testing the revised prototype, they realized that viewing a video script on mobile was overwhelming. They felt this realization could have only occurred on MobileMaker: “Walking around and trying it on mobile gave me a much better intuition on what I liked and disliked… for example, generating a full script for the video ideas on the go is overwhelming and too much text. I would have probably said I wanted that feature on desktop without realizing this. That realization is much more likely to happen if you’re actually using it, as opposed to being asked to think about it.” Beyond this, NL revision also empowered users to make subsequent, “cascading” revisions (or cascading feedback based on a revision). In the example above, when P8 added the dropdown menu for selection of “vibe” or “object”, they found that the model actually had difficulty assessing vibe from one image alone (described in sec VI-A4).
@@ VI-B3 Revising helped shorten the test-feedback loop in MobileMaker
Revising in MobileMaker helped participants feel like they were actively and iteratively improving the prototype. While experiencing Img2Playlist, P2 added a dropdown menu for genre to control the types of songs output by the application and then added a description to accompany each song to explain their relevance to the input image. They felt that these additions tangibly improved the base application by introducing greater user control of the playlist and tightening the connection between the input image and output playlist: “In the mobile prototype, I felt like we were stepping up the stairs to get to a better and better, more refined version of the north-star version of the application.” In the baseline, even though their feedback was used to alter the prompt in the editor, P2 felt they were giving feedback on the prompt’s output, but the application as a whole was not improving. Echoing these thoughts, P16 described how experiencing their suggested revisions felt more active and fulfilling: “You can complete this loop of giving feedback and seeing it in action. This experience is definitely much better, rather than describing your frustrations, and hopefully getting them fixed.” Participants’ revisions in MobileMaker would build on each other, which helped them feel (1) more active in the design process and (2) that they were tangibly improving the application.
VI-B4 Participants felt restricted to give feedback and revisions enabled by MobileMaker
While participants enjoyed immediately experiencing their feature suggestions, many also felt restricted by the revisions that were possible in MobileMaker. P6 explained, “With [MobileMaker ], you’re a bit limited by what you can add to the UI. And whatever revision you add will be a bit more permanent in the mobile prototype.” With MobileMaker, participants tended to focus on feedback that could lead to new revisions they could experience, e.g., adding a dropdown or a text input to steer the output playlist, potentially at the expense of providing feedback that could not be immediately rendered as a revision (e.g., thoughts on alternative user journeys or workflows). Overall, a potential drawback of giving users the ability to revise their prototype is that they focus on feedback that can lead to revisions, instead of critically thinking about the application at a high-level.
@@ VI-B5 Challenges in implementing revisions
While we intended for the participants to be able to independently update the prototypes, in practice, the facilitators occasionally needed to step in to help. In general, the LLM prompt powering revisions worked best when users (1) input a single revision and (2) clearly specified that revision. On a few occasions, participants entered revision requests with multiple changes (e.g., add two separate dropdowns and remove a text input) and the prompt failed to produce all of the specified changes. This type of error was mostly mitigated by educating users on the feature during our short demonstration. One more challenging issue was clearly delineating requests. While P13 was testing Img2VideoIdeas, they wanted the application to take in multiple images that would be used together to inform three video ideas. They wrote: “I want three photos to be considered together,” but the model interpreted this to mean creating a separate video idea for each image. If P13 had better specified their request, it likely would have been implemented correctly; however, users often wanted to provide a quick description and have the model correctly extrapolate their intent.
@@ VII Design Reflection Exercise
In addition to the exploratory user study, we also conducted an initial probe on how designers might interact with users’ prototype revisions. From the MobileMaker condition of the previous study, we presented a dashboard of participants’ suggested revisions, as well as their revised prototypes, to the two professional designers who had initially prototyped Img2Playlist and Img2VideoIdeas. Designers spent approximately half an hour perusing the dashboard and applying the suggested revisions, and were asked to think aloud.
A designer remarked that, relative to traditional forms of user feedback, experiencing users’ dynamically-runnable revised prototypes was “very powerful and a more convincing form of feedback.” They thought it made user feedback more tangible and compelling compared to vague requests such as “I want [feature request],” which are easier for designers to neglect or dismiss in practice. Beyond this, both designers found it crucial to have access to the in-the-wild context embedded within those revised prototypes. In MobileMaker, designers could run the revised prototype on the actual inputs experienced by the user (e.g., picture of pool table), and see the outputs the user saw. Those test cases also served as tangible evidence for suggested revisions. For example, the designers were initially perplexed by a sudden request to change the original Img2Playlist feature to “video2playlist.” After noticing that the participant had previously inputted a series of photos depicting different scenes in a game room, they hypothesized that the participant preferred video over still images to better capture the dynamic context and ambiance. In the future, designers would like to see more of the “user journey” and reasoning behind the proposed revisions, so they could better understand users’ motivations while running the prototypes.
Designers wished that there was an easier way to synthesize the different threads of feedback that participants provided in MobileMaker so that they could be more effectively utilized. One proposed solution was for MobileMaker to automatically organize the suggested revisions into distinct themes, such as “providing rationale for song choices in the output,” “using video instead of image as input,” or “add more UI controls to influence the type of playlist,” as well as statistics on their frequency. As one designer explained, “it’s not just one person reading through [the feedback] but a whole team that needs to be convinced that this needs to happen in this way,” so it would be important to have sufficient evidence to help prioritize and facilitate more targeted design adjustments.
@@ VIII-A Agile AI-Prototyping Practices in the Age of LLMs
Short design iteration loops coupled with frequent testing have long been advocated for in the design and software development communities. This paper expands on the question of what might be possible during the early phases of AI design if (1) designers can quickly push functional AI prototypes out into the wild and (2) testers can contribute both feedback and updated functional prototypes while in the field (Fig. 6). In light of this, we may witness a parallel shift in the roles and responsibilities of designers, user researchers, developers, and end-users. Designers and developers might begin to borrow practices from user research (e.g., in-the-field user observations); likewise, user researchers may take on more designerly and creative roles by troubleshooting or refining user-proposed revisions directly in the field. Overall, in situ prototyping could introduce a blurring of boundaries between these traditional roles, broadening participation of AI development to include more diverse professions and end-users. While this may inevitably introduce new frictions (e.g., increased coordination costs of disambiguating roles, potential suboptimal designs by end-users [60], etc.), future work could also capitalize on the new opportunities brought on by this blurring of roles, e.g., designers and developers more deeply considering off-the-desktop user needs; UX researchers taking a more active role in implementing user feedback; and end-user communities personalizing and configuring AI to their own needs.
Figure 6: Traditional (left) vs. in situ (right) prototyping cycles.
@@ VIII-E Limitations and Future Work
During the MobileMaker portion of the study, participants mostly explored areas inside or adjacent to their office buildings. To enrich results, users could ideally use the applications in a variety of natural settings, such as near their homes or at favorite spots in the city. However, for practicality and to observe users as they experienced the MobileMaker prototypes in person, we kept the locations standard across participants. While MobileMaker’s feature set is not exhaustive, it enabled us to probe the value and opportunities of in situ AI prototyping and testing in our exploratory study. In the future, MobileMaker can be expanded to support building prototypes with multi-screen workflows and a wider array of UI inputs and outputs.
@@ IX Conclusion
To explore the opportunities for in situ AI prototyping and testing, we developed MobileMaker, with which designers can (1) rapidly create mobile AI prototypes and (2) enable users to revise the prototype’s design through natural language in real-time. In our exploratory user study, we found that with MobileMaker prototypes, users can authentically experience the envisioned application, which in turn helped them evaluate if the AI’s outputs are mobile appropriate, serendipitously experiment with image edge cases, discover discrepancies between their interpretation of the problem and the AI’s interpretation, and more critically evaluate the AI’s outputs via in-the-wild contextual cues. In addition, NL revision enabled users to provide better guardrails to the AI in response to in-the-wild discrepancies, critically evaluate their own feedback in situ, and shorten the test-feedback loop. Two potential avenues for future work include further supporting (1) users with providing feedback on in situ medium fidelity prototypes, as well as (2) designers with collecting and deriving insights from this feedback at scale.

### v2

@@ VI-B2 NL revise enabled testers to critically evaluate their own feedback in situ
With MobileMaker’s NL revision feature, testers could immediately “live” and experience their own feedback within the updated prototype, which often helped them rethink their initial feedback. P4 initially revised Img2VideoIdeas to “Add a script for each video idea”, but after testing the revised prototype, they realized that viewing a video script on mobile was overwhelming. They felt this realization could have only occurred on MobileMaker: “Walking around and trying it on mobile gave me a much better intuition on what I liked and disliked… for example, generating a full script for the video ideas on the go is overwhelming and too much text. I would have probably said I wanted that feature on desktop without realizing this. That realization is much more likely to happen if you’re actually using it, as opposed to being asked to think about it.” Beyond this, NL revision also empowered testers to make subsequent, “cascading” revisions (or cascading feedback based on a revision). In the example above, when P8 added the dropdown menu for selection of “vibe” or “object”, they found that the model actually had difficulty assessing vibe from one image alone (described in sec VI-A4).
@@ VI-B3 Revising helped shorten the test-feedback loop in MobileMaker
Revising in MobileMaker helped testers feel like they were actively and iteratively improving the prototype. While experiencing Img2Playlist, P2 added a dropdown menu for genre to control the types of songs output by the application and then added a description to accompany each song to explain their relevance to the input image. They felt that these additions tangibly improved the base application by introducing greater user control of the playlist and tightening the connection between the input image and output playlist: “In the mobile prototype, I felt like we were stepping up the stairs to get to a better and better, more refined version of the north-star version of the application.” In the baseline, even though their feedback was used to alter the prompt in the editor, P2 felt they were giving feedback on the prompt’s output, but the application as a whole was not improving. Echoing these thoughts, P16 described how experiencing their suggested revisions felt more active and fulfilling: “You can complete this loop of giving feedback and seeing it in action. This experience is definitely much better, rather than describing your frustrations, and hopefully getting them fixed.” Testers’ revisions in MobileMaker would build on each other, which helped them feel (1) more active in the design process and (2) that they were tangibly improving the application.
VI-B4 Testers felt restricted to give feedback and revisions enabled by MobileMaker
While testers enjoyed immediately experiencing their feature suggestions, many also felt restricted by the revisions that were possible in MobileMaker. P6 explained, “With [MobileMaker ], you’re a bit limited by what you can add to the UI. And whatever revision you add will be a bit more permanent in the mobile prototype.” With MobileMaker, testers tended to focus on feedback that could lead to new revisions they could experience, e.g., adding a dropdown or a text input to steer the output playlist, potentially at the expense of providing feedback that could not be immediately rendered as a revision (e.g., thoughts on alternative user journeys or workflows). Meanwhile, in the baseline condition, testers often ideated more “out there” feature requests and entirely different functionalities with Figma. For example, P14 suggested that Img2VideoIdeas could be embedded into existing social media platforms and altered to also support caption generation for video posts. This higher level, more “out there” feedback elicited through Figma might be more useful to designers during early stage ideation, whereas the more granular and grounded feedback enabled by revising might be more useful further along the design process, when designers are refining their application’s functionality. Overall, because they elicit quite different feedback, Figma and MobileMaker might be employed at different parts of the design process, depending on the needs of the designer.
@@ VI-B5 Challenges in implementing revisions
While we intended for the testers to be able to independently update the prototypes, in practice, the facilitators occasionally needed to step in to help. In general, the LLM prompt powering revisions worked best when testers (1) input a single revision and (2) clearly specified that revision. For example, revision requests that added a single input like “add a dropdown for a music genre” or made a simple, clear adjustment to the model’s output, e.g. “give me 5 songs instead of 10” tended to work well immediately. However, on a few occasions, testers entered revision requests with multiple changes (e.g., add two separate dropdowns and remove a text input) and the prompt failed to produce all of the specified changes. This type of error was mostly mitigated by educating testers on the feature during our short demonstration. One more challenging issue was clearly delineating requests. While P13 was testing Img2VideoIdeas, they wanted the application to take in multiple images that would be used together to inform three video ideas. They wrote: “I want three photos to be considered together,” but the model interpreted this to mean creating a separate video idea for each image. If P13 had better specified their request, it likely would have been implemented correctly; however, testers often wanted to provide a quick description and have the model correctly extrapolate their intent.
@@ VII Design Reflection Exercise
In addition, we conducted an initial probe on how designers might interact with testers’ prototype revisions. From the MobileMaker condition of the previous study, we presented a dashboard of participants’ suggested revisions, as well as their revised prototypes, to the two designers who had initially prototyped Img2Playlist and Img2VideoIdeas. Designers spent around 30 minutes perusing the dashboard and applying the suggested revisions while thinking aloud.
D1 remarked that, relative to traditional forms of user feedback, experiencing users’ dynamically-runnable revised prototypes was “very powerful and a more convincing form of feedback.” They thought it made user feedback more tangible and compelling compared to vague requests such as “I want [feature request],” which are easier for designers to neglect or dismiss in practice. Beyond this, both designers found it crucial to have access to the in-the-wild context embedded within those revised prototypes. In MobileMaker, designers could run the revised prototype on the actual inputs experienced by the user (e.g., picture of pool table), and see the outputs the user saw. Those test cases also served as tangible evidence for suggested revisions. For example, the designers were initially perplexed by a sudden request to change the original Img2Playlist feature to “video2playlist.” After noticing that the participant had previously inputted a series of photos depicting different scenes in a game room, they hypothesized that the participant preferred video over still images to better capture the dynamic context and ambiance. In the future, designers would like to see more of the “user journey” and reasoning behind the proposed revisions, so they could better understand users’ motivations while running the prototypes.
Designers wished that there was an easier way to synthesize the different threads of feedback that participants provided in MobileMaker so that they could be more effectively utilized. One proposed solution by D2 was for MobileMaker to automatically organize the suggested revisions into distinct themes, such as “providing rationale for song choices in the output,” “using video instead of image as input,” or “add more UI controls to influence the type of playlist,” as well as statistics on their frequency. As one designer explained, “it’s not just one person reading through [the feedback] but a whole team that needs to be convinced that this needs to happen in this way,” so it would be important to have sufficient evidence to help prioritize and facilitate more targeted design adjustments.
@@ VIII-A Agile AI-Prototyping Practices in the Age of LLMs
Short design iteration loops coupled with frequent testing have long been advocated for in the design and software development communities. This paper expands on the question of what might be possible during the early phases of AI design if (1) designers can quickly push functional AI prototypes out into the wild and (2) testers can contribute both feedback and updated functional prototypes while in the field. In light of this, we may witness a parallel shift in the roles and responsibilities of designers, user researchers, developers, and end-users. Designers and developers might begin to borrow practices from user research (e.g., in-the-field user observations); likewise, user researchers may take on more designerly and creative roles by troubleshooting or refining user-proposed revisions directly in the field. Overall, in situ prototyping could introduce a blurring of boundaries between these traditional roles, broadening participation of AI development to include more diverse professions and end-users. While this may inevitably introduce new frictions (e.g., increased coordination costs of disambiguating roles, potential suboptimal designs by end-users [61], etc.), future work could also capitalize on the new opportunities brought on by this blurring of roles, e.g., designers and developers more deeply considering off-the-desktop user needs; UX researchers taking a more active role in implementing user feedback; and end-user communities personalizing and configuring AI to their own needs.
@@ VIII-E Limitations and Future Work
During the MobileMaker portion of the study, participants mostly explored areas inside or adjacent to their office buildings. To enrich results, participants could ideally use the applications in a variety of natural settings, such as near their homes or at favorite spots in the city. However, for practicality and to observe participants as they experienced the MobileMaker prototypes in person, we kept the locations standard across participants. Also, future work can examine how MobileMaker impacts tester feedback on a wider variety of applications. Finally, while MobileMaker’s feature set is not exhaustive, it enabled us to probe the value and opportunities of in situ AI prototyping and testing in our exploratory study. In the future, MobileMaker can be expanded to support building prototypes with multi-screen, multi-LLM calls workflows and a wider array of UI inputs and outputs.
@@ IX Conclusion
To explore the opportunities for in situ AI prototyping and testing, we developed MobileMaker, with which designers can (1) rapidly create mobile AI prototypes and (2) enable testers to authentically experience the prototype in the wild and revise it in real-time. In our exploratory study, we found that by testing in the wild, testers were able to serendipitously experiment with image edge cases. In addition, NL revision enabled testers to critically evaluate their own feedback but also potentially limited the feedback they gave to what was actionable in MobileMaker. Future work might explore further supporting (1) users with providing feedback on in situ medium fidelity prototypes, as well as (2) designers with deriving insights from this feedback at scale.